OpenStudy (anonymous):

A circle has a radius of 2.5 centimeters and a central angle AOB that measures 90°. What is the area of sector AOB? Use 3.14 for pi and round your answer to the nearest tenth. A. 2.9 cm2 B. 3.9 cm2 C. 4.9 cm2 D. 19.6 cm2

OpenStudy (noahbred):

is this for online schooling or regular schooling

OpenStudy (dayakar):

Do u know formula for area of a sector

OpenStudy (dayakar):

Radius (r)=2.5 cm Angle(x)= 90 Area of the sector = \[\frac{ x }{ 360 }×\pi×r ^{2}\]

OpenStudy (dayakar):

Plug in x, r, values and get the area of the sector

OpenStudy (anonymous):

3.14 (2.5)^2 (1/4) = 4.9 cm^2